OpenStudy (anonymous):

Help! The midpoint of CD is E(–1, 0). One endpoint is C(5, 2). What are the coordinates of the other endpoint? Explain the process...

OpenStudy (anonymous):

I use midpoint formula x=1/2 (x1 + x2) Assuming x1=5 And x=-1 (midpoint) -1=1/2 (5+ x2) -2=5+x2 x2= -7 y=0 Y1=2 Y=1/2 (y1+y2) 0=1/2(2+y2) 0=2+y2 y2=-2 hence, coordinates are (-7,-2)